# Unveiling the Ignorance of MLLMs: Seeing Clearly, Answering Incorrectly

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have displayed remarkable performance in multi-modal tasks, particularly in visual comprehension. However, we reveal that MLLMs often generate incorrect answers even when they understand the visual content. To this end, we manually construct a benchmark with 12 categories and design evaluation metrics that assess the degree of error in MLLM responses even when the visual content is seemingly understood. Based on this benchmark, we test 15 leading MLLMs and analyze the distribution of attention maps and logits of some MLLMs. Our investigation identifies two primary issues: 1) most instruction tuning datasets predominantly feature questions that “directly” relate to the visual content, leading to a bias in MLLMs’ responses to other indirect questions, and 2) MLLMs’ attention to visual tokens is notably lower than to system and question tokens. We further observe that attention scores between questions and visual tokens as well as the model’s confidence in the answers are lower in response to misleading questions than to straightforward ones. To address the first challenge, we introduce a paired positive and negative data construction pipeline to diversify the dataset. For the second challenge, we propose to enhance the model’s focus on visual content during decoding by refining the text and visual prompt. For the text prompt, we propose a content guided refinement strategy that performs preliminary visual content analysis to generate structured information before answering the question. Additionally, we employ a visual attention refinement strategy that highlights question-relevant visual tokens to increase the model’s attention to visual content that aligns with the question. Extensive experiments demonstrate that these challenges can be significantly mitigated with our proposed dataset and techniques. The benchmark, training set, and code will be available.
@@
Due to the space constraints of the main manuscript, this Supplementary Material provides a comprehensive presentation of additional details and analyses. The content is structured as follows: (1) a detailed overview of the MMVU test set, including its construction criteria and a preliminary evaluation conducted on a subset of the test set; (2) findings from initial exploratory experiments to assess the test set’s effectiveness using a subset of the MMVU dataset; (3) an examination of the MLLM’s performance under various data combination strategies, prompt versions, and subsets of MMVU-Train of differing sizes, accompanied by an in-depth discussion on the construction of training prompts and their impact on model performance; (4) the training of an MLLM using the MMVU-Train dataset, validated on both a subset of the MMVU test set and other general datasets; (5) representative examples of paired positive and negative samples; and (6) the specific content of the ”Instruction” prompt employed during model testing.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have garnered significant attention due to their impressive capabilities in addressing diverse multimodal tasks, particularly in visual comprehension and reasoning [72, 4, 26]. Despite these strengths, MLLMs sometimes produce responses that deviate from the image content, resulting in what is termed “hallucinations”. Previous studies [18] like MADBench [67], MME [21], POPE [47], ConBench [83], and NaturalBench [41] introduce datasets for evaluating the accuracy or consistency of MLLMs’ responses under image or text perturbations. However, they do not distinguish between the causes of incorrect responses. We observe that incorrect responses may not just stem from the MLLMs not understanding the visual content. Instead, we find that it is quite prevalent for an MLLM to understand the content of the image but yet still generate incorrect answers, as illustrated in Fig. 1 (a)). Our findings suggest that subtle phrasing in questions can introduce biases that hinder MLLMs from leveraging their understanding accurately. This raises two critical questions for our research: 1) Why does the model understand the visual content but provide incorrect answers? 2) How can we correct this error-prone behavior of MLLMs?
@@ 2 Related Works
Multimodal Large Language Models (MLLMs). Building on the success of LLMs, recent research has increasingly focused on MLLMs [2, 72]. This shift aims to achieve better cross-modality understanding and reasoning. Currently, a significant amount of research is exploring various aspects of MLLMs, including the design of structures [55, 48], data construction [37, 49, 85], training strategies [62, 33, 58], and lightweight MLLMs [26, 88]. For a comprehensive review of these topics, readers are encouraged to investigate surveys [82, 32, 7] dedicated to several areas.
@@ 3.2 MMVU Benchmark Design
As illustrated in Fig. 3, we design paired positive and negative questions for each sample to ensure a comprehensive evaluation. Each question offers four brief options, with only one correct answer. In total, we manually construct 1,786 questions (893 positive and 893 negative) across three levels: character, attribute, and context. Character-level questions focus on elements such as characters or numbers, attribute-level questions assess properties like color, texture, and quantity, and context-level questions explore higher-level concepts such as emotions, culture, and common sense. The positive questions are designed to evaluate the model’s comprehension abilities, while the negative questions test its resistance to interference. For example, misleading character-level questions alter elements like characters or numbers, attribute-level questions introduce property confusion, and context-level questions challenge the model with complex concepts, thoroughly assessing its robustness against misleading information.
Figure 4: Please refer to Section 3.4 for details on the following calculations. (a) Statistical results of attention scores between the answer tokens and the system, visual, and question tokens. It is seen that the answer tokens pay the least attention to the visual tokens in general. (b) The ratio of the question tokens’ attention to the system and visual tokens in negative samples versus positive samples. Negative questions appear to pay less attention than positive questions to visual tokens. (c) Procedure for calculating the ratio of output probabilities for negative and positive samples. (d) Comparison of the ratio of output probabilities for negative and positive samples across different MLLMs. Interestingly, it seems that a lower output probability correlates with lower attention between the question and visual tokens in (b).
@@ 3.3 Evaluation metrics
We develop metrics grounded in Bayesian conditional probabilities to evaluate whether an MLLM comprehends visual content and its robustness against misdirection. Specifically, we categorize the results into four groups: 1) Understanding and Robustness (UR), 2) Understanding but Fragility (UF), 3) Not Understanding and Rigorous (NR), and 4) Not Understanding and Fragility (NF). To quantify the effect of negative questions in misleading the model despite its comprehension of visual content, we introduce the “Misleading Rate (MR)”. Additionally, we introduce the “Robustness Accuracy (RA)” to evaluate the understanding capability of MLLMs. The formulation is as follows:
@@ 3.5 MMVU Training Set Construction Pipeline
Information extraction. Previous efforts that generate instruction-tuning data for multimodal conversations with GPT-4o primarily fall into two categories: direct generation and annotation-driven generation. Direct generation approaches, such as ALLaVA [9], rely entirely on the basic ability of GPT-4o to synthesize samples. In contrast, annotation-driven generation methods, such as LRV [52], utilize existing image annotations (e.g. object bounding boxes and textual descriptions) as additional information for text-only GPT-4 to generate conversational samples. However, these works fail to make full use of fine-grained details. On the contrary, our data pipeline adopts a different scheme by implicitly and comprehensively extracting the detailed information directly from the image. This includes 1) text or numbers (if present), 2) objects and people, 3) object attributes (e.g., colors, textures, locations) and human characteristics (e.g., postures, dresses), 4) interrelationships between people and objects, 5) events or activities, and 6) the overall feeling and understanding evoked by the image. The detailed prompts are in Supplementary Material.
@@ 4.1 Analysis of MMVU Benchmark
Table 4: Ablation study on understanding-based prompting on the MMVU test set. Blue and Red text indicate the extent of performance improvement and decrease, respectively.
@@ 4.3 Analysis of MMVU Training Set
Due to page limitations, more quantitative, visual results, and discussion, are provided in the Supplementary Material.
Experiment results on the MMVU benchmark. To evaluate our proposed dataset, we fine-tune several state-of-the-art MLLMs, including LLaVA-OneVision-7B [54], InternVL2-8B [14], VILA1.5-13B [58], and Phi3.5-Vision-Instruct [1], using MMVU training set (Section 3.5) and assess their performance on the MMVU benchmark. As shown in Tab. 2, models fine-tuned with our dataset consistently outperformed their baselines across most metrics. Notably, LLaVA-OneVision-7B + MMVU-Train demonstrated improvements in character/number recognition (68.29% vs. 62.20%), numerical understanding (55.56% vs. 41.98%), and activity recognition (80.00% vs. 75.56%). These results indicate that the paired positive-negative data significantly reduces biased responses.
@@ 5 Conclusion
Overview of Supplementary Material
[…]
Supplementary Materials Contents
Appendix A MMVU Test set Details
@@
A.1 Category Definition and Negative Question Design
@@
A.2 Preliminary Evaluation Details.
@@
Appendix B Discussion about the Effectiveness of the MMVU Test Set
@@
Appendix C MMVU-Train Data Construction Preliminary Experiment
C.1 Notes
@@
C.2 Preliminary Testing Data Distribution.
We have collected images that matched 12 subcategory themes from the web and some test sets, such as Flickr-30k [78], and facilitated setting misleading prompts. We meticulously design questions and standard answers based solely on the image information without relying on the original captions. The distribution of the subset of the MMVU test set is shown in Fig. 2. It categorizes questions into three main types: ”context” (120 questions), “character” (40 questions), and “attribute” (140 questions). The “context” category includes subtypes such as relationships (22), activity (24), expertise (31), concrete (21), and abstract (22) questions. The “character” category comprises character/number questions (40). The “attribute” category encompasses presence (22), color/texture (24), number (24), shape (24), posture (24), and position (22) questions. This distribution highlights a comprehensive and balanced approach, ensuring a well-rounded benchmark with detailed subcomponents across various dimensions of question types.
C.3 Training Details
@@
C.4 Discussion 1: Can We Simply Introduce Negative Samples to Train MLLMs?
@@
1. If there is no text/number recognized or you do not make sure the text/number is correct, do not ask this type of question. If text is present, inquire about modifications to characters or numbers: Generate questions involving minor alterations, such as reversing the order of digits (modifying a number) or replacing a character. 2. If characters or numbers are present in the image, modify one letter of a word or one digit of a number, then ask if the modified character or number exists, and add a distractor sentence to make the question more challenging. Otherwise, do not ask this type of question. 3. Ask if certain objects exist: Replace the nouns of objects in the image with visually similar but nonexistent objects, then ask if the model identifies the object. 4. Ask if objects exist and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. 5. If objects have attributes such as names of people, colors of objects, counts of objects, or positions, ask modified questions about object attributes and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. 6. Ask global understanding questions: Understand the broader context, culture, emotions, and knowledge depicted in the image. Pose questions that differ slightly from true understanding but are very similar. 7. Ask global understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. 8. Ask local understanding questions: Based on the extracted relationships between objects, between people and objects, or between people, actions, or relative positions, pose local understanding challenges (negative samples). 9. Ask local understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. Requirements: 1. Generate 7-10 questions. 2. Ask directly, without adding any information such as ”in the conversation,” ”mentioned in the conversation,” or ”in the image.” 3. Answer directly, with explanation, succinctly, with responses not exceeding 30 characters. 4. Questions must be real and answerable. Answers must be correct, definitely,and not speculative. 5. Do not generate questions that the answer is ”uncertain”. 6. The format of the generated questions and answers is: ’conversations’: [’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, …, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’] STEP-3: Output JSON sample
@@
1. Character: (a) If no text/number is recognized or the text/number recognized is not sure, do not ask questions about text/number. If text exists, inquire about modifications to characters or numbers: Generate questions involving minor alterations, such as reversing the order of digits (modifying a number) or replacing a character. (b) If there are characters or numbers in the image, modify one letter or one digit, then ask if the modified character or number exists, and add a distractor sentence to make the question more challenging. Otherwise, do not ask this type of question. 2. Semantic: (a) Ask if certain objects exist: Replace the nouns of objects in the image with visually similar but nonexistent objects, then ask if the model identifies the object. (b) Ask if objects exist and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. (c) If objects have attributes such as names of people, colors of objects, counts of objects, or positions, ask modified questions about object attributes and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. 3. Understanding: (a) Ask global understanding questions: Understand the broader context, culture, emotions, and knowledge depicted in the image. Pose questions that differ slightly from true understanding but are very similar. (b) Ask global understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. (c) Ask local understanding questions: Based on the extracted relationships between objects, between people and objects, or between people, actions, or relative positions, pose local understanding challenges (negative samples). (d) Ask local understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. Requirements: 1. Generate 2 character-type questions, 4 semantic-type questions, and 2 understanding-type questions. 2. Ask directly, without adding any information such as ”in the conversation,” ”mentioned in the conversation,” or ”in the image.” 3. Answer directly, with explanation, succinctly, with responses not exceeding 30 characters. 4. Questions must be real and answerable. Answers must be correct, definitely,and not speculative. 5. Do not generate questions that the answer is ”uncertain”. 6. The format of the generated questions and answers is: ’conversations’: [’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, …, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’] STEP-3: Output JSON sample
@@
C.4.1 Training and Results of the Version 0.
@@
C.4.2 Training and Results of the Version 1.
@@
C.5 Discussion 2: What Constitutes an Effective Pair of Positive and Negative Samples for Training?
Having established the benefit of incorporating both positive and negative samples during training, we now delve into effective construction methods for such pairs. We investigate two distinct approaches: 1) Explicit Positive/negative Prompting. This method involves directly prompting the model to generate both positive and negative samples, with answers explicitly labeled as ”yes” or ”no.” This approach offers straightforward control over the positive/negative balance in the training data (version 2). Confounding Answer Construction. This approach focuses on crafting questions that naturally lend themselves to multiple plausible answers, without explicitly forcing a ”yes/no” response. This strategy aims to create a more nuanced and challenging training dataset, potentially enhancing the model’s ability to discern subtle differences in meaning and context (version 3). The GPT-4V prompts for version 2 and version 3 are shown in Fig. 9 and Fig. 10, respectively. To evaluate the effectiveness of these two versions of prompts, we generated 24k positive-negative sample pairs using each version’s prompt. We then incorporated these datasets into our training pipeline, maintaining consistent experimental settings with previous trials. Fig. 11 presents the training loss curves observed for both approaches, while Tab. 10 and Tab. 11 detail the corresponding performance results across various benchmarks. Our experimental results demonstrate that utilizing prompt version 3 to generate paired positive and negative samples yielded the strongest overall performance. Consequently, we adopted version 3 as our preferred method for generating training data in the main experiments.

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) have displayed remarkable performance in multi-modal tasks, particularly in visual comprehension. However, we reveal that MLLMs often generate incorrect answers even when they understand the visual content. To this end, we manually construct a benchmark with 12 categories and design evaluation metrics that assess the degree of error in MLLM responses even when the visual content is seemingly understood. Based on this benchmark, we test 15 leading MLLMs and analyze the distribution of attention maps and logits of some MLLMs. Our investigation identifies two primary issues: 1) most instruction tuning datasets predominantly feature questions that “directly" relate to the visual content, leading to a bias in MLLMs’ responses to other indirect questions, and 2) MLLMs’ attention to visual tokens is notably lower than to system and question tokens. We further observe that attention scores between questions and visual tokens as well as the model’s confidence in the answers are lower in response to misleading questions than to straightforward ones. To address the first challenge, we introduce a paired positive and negative data construction pipeline to diversify the dataset. For the second challenge, we propose to enhance the model’s focus on visual content during decoding by refining the text and visual prompt. For the text prompt, we propose a content guided refinement strategy that performs preliminary visual content analysis to generate structured information before answering the question. Additionally, we employ a visual attention refinement strategy that highlights question-relevant visual tokens to increase the model’s attention to visual content that aligns with the question. Extensive experiments demonstrate that these challenges can be significantly mitigated with our proposed dataset and techniques.
@@
Due to the space constraints of the main manuscript, this Supplementary Material provides a comprehensive presentation of additional details and analyses. The content is structured as follows: (1) a detailed overview of the MMVU test set, including its construction criteria and a preliminary evaluation conducted on a subset of the test set; (2) findings from initial exploratory experiments to assess the test set’s effectiveness using a subset of the MMVU dataset; (3) an examination of the MLLM’s performance under various data combination strategies, prompt versions, and subsets of MMVU-Train of differing sizes, accompanied by an in-depth discussion on the construction of training prompts and their impact on model performance; (4) the training of an MLLM using the MMVU-Train dataset, validated on both a subset of the MMVU test set and other general datasets; (5) representative examples of paired positive and negative samples; and (6) the specific content of the "Instruction" prompt employed during model testing.
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have garnered significant attention due to their impressive capabilities in addressing diverse multimodal tasks, particularly in visual comprehension and reasoning [77, 4, 27]. Despite these strengths, MLLMs sometimes produce responses that deviate from the image content, resulting in what is termed “hallucinations". Previous studies [19] like MADBench [69], MME [22], POPE [48], ConBench [88], and NaturalBench [42] introduce datasets for evaluating the accuracy or consistency of MLLMs’ responses under image or text perturbations. However, they do not distinguish between the causes of incorrect responses. We observe that incorrect responses may not just stem from the MLLMs not understanding the visual content. Instead, we find that it is quite prevalent for an MLLM to understand the content of the image but yet still generate incorrect answers, as illustrated in Fig. 1 (a)). Our findings suggest that subtle phrasing in questions can introduce biases that hinder MLLMs from leveraging their understanding accurately. This raises two critical questions for our research: 1) Why does the model understand the visual content but provide incorrect answers? 2) How can we correct this error-prone behavior of MLLMs?
@@ 2 Related Works
Multimodal Large Language Models (MLLMs). Building on the success of LLMs, recent research has increasingly focused on MLLMs [2, 77, 75, 74, 10, 62]. This shift aims to achieve better cross-modality understanding and reasoning. Currently, a significant amount of research is exploring various aspects of MLLMs, including structure design [56, 49, 76], data construction [38, 50, 90], training strategies [64, 34, 59], and lightweight MLLMs [27, 93]. For a comprehensive review of these topics, readers are encouraged to investigate surveys [87, 33, 7].
@@ 3.2 MMVU Benchmark Design
As illustrated in Fig. 3, we design paired positive and negative questions to ensure a comprehensive evaluation. Each question offers four brief options, with only one correct answer. In total, we manually construct 1,786 questions (893 positive and 893 negative) across three levels: character, attribute, and context. Character-level questions focus on elements such as characters or numbers, attribute-level questions assess properties like color, texture, and quantity, and context-level questions explore higher-level concepts such as emotions, culture, and common sense. The positive questions are designed to evaluate the model’s comprehension abilities, while the negative questions test its resistance to interference. For example, misleading character-level questions alter elements like characters or numbers, attribute-level questions introduce property confusion, and context-level questions challenge the model with complex concepts, assessing its robustness against misleading information.
Figure 4: Please refer to Sec. 3.4 for details on the following calculations. (a) Statistical results of attention scores between the answer tokens and the system, visual, and question tokens. It is seen that the answer tokens pay the least attention to the visual tokens in general. (b) The ratio of the question tokens’ attention to the system and visual tokens in negative samples versus positive samples. Negative questions appear to pay less attention than positive questions to visual tokens. (c) Procedure for calculating the ratio of output probabilities for negative and positive samples. (d) Comparison of the ratio of output probabilities for negative and positive samples across different MLLMs. Interestingly, it seems that a lower output probability correlates with lower attention between the question and visual tokens in (b).
@@ 3.3 Evaluation metrics
We develop metrics grounded in Bayesian conditional probabilities to evaluate whether an MLLM comprehends visual content and its robustness against misdirection. Specifically, we categorize the results into four groups: 1) Understanding and Robustness (UR): The model accurately comprehends the visual content and remains resilient when exposed to negative questions. 2) Understanding but Fragility (UF): The model successfully understands the visual content but is vulnerable to errors when faced with negative questions. 3) Not Understanding and Rigorous (NR): The model fails to grasp the visual content but still provides correct answers to negative questions. 4) Not Understanding and Fragility (NF): The model neither comprehends the visual content nor answers the negative questions correctly. To quantify the effect of negative questions in misleading the model despite its comprehension of visual content, we introduce the “Misleading Rate (MR)”. Additionally, we introduce the “Robustness Accuracy (RA)” to evaluate the understanding capability of MLLMs. The formulation is as follows:
@@ 3.5 MMVU Training Set Construction Pipeline
Information extraction. Previous efforts that generate instruction-tuning data for multimodal conversations with GPT-4o primarily fall into two categories: direct generation and annotation-driven generation. Direct generation approaches, such as ALLaVA [9], rely entirely on the basic ability of GPT-4o to synthesize samples. In contrast, annotation-driven generation methods, such as LRV [53], utilize existing image annotations (e.g. object bounding boxes and textual descriptions) as additional information for text-only GPT-4 to generate conversational samples. However, these works fail to make full use of fine-grained details. On the contrary, our data pipeline adopts a different scheme by implicitly and comprehensively extracting the detailed information directly from the image. This includes 1) text or numbers (if present), 2) objects and people, 3) object attributes (e.g., colors, textures, locations) and human characteristics (e.g., postures, dresses), 4) interrelationships between people and objects, 5) events or activities, and 6) the overall feeling and understanding evoked by the image.
@@ 4.1 Analysis of MMVU Benchmark
Table 4: Ablation study on understanding-based prompting in MMVU. Blue and Red indicate performance gains and drops, respectively.
@@ 4.3 Analysis of MMVU Training Set
Experiment results on the MMVU benchmark. To evaluate our proposed dataset, we fine-tune several state-of-the-art MLLMs, including LLaVA-OneVision-7B [55], InternVL2-8B [15], VILA1.5-13B [59], and Phi3.5-Vision-Instruct [1], using MMVU training set (Sec. 3.5) and assess their performance on the MMVU benchmark. Weight merging is utilized in the experiment. As shown in Tab. 2, models fine-tuned with our dataset consistently outperformed their baselines across most metrics. Notably, LLaVA-OneVision-7B + MMVU-Train demonstrated improvements in character/number recognition (68.29% vs. 62.20%), numerical understanding (55.56% vs. 41.98%), and activity recognition (80.00% vs. 75.56%). These results indicate that the paired positive-negative data can reduce biased responses.
@@ 5 Conclusion
6 MMVU Test set Details
@@
6.1 Category Definition and Negative Question Design
@@
6.2 Preliminary Evaluation Details.
@@
7 Discussion about the Effectiveness of the MMVU Test Set
@@
8 MMVU-Train Data Construction Preliminary Experiment
8.1 Notes
@@
8.2 Preliminary Testing Data Distribution.
We have collected images that matched 12 subcategory themes from the web and some test sets, such as Flickr-30k [83], and facilitated setting misleading prompts. We meticulously design questions and standard answers based solely on the image information without relying on the original captions. The distribution of the subset of the MMVU test set is shown in Fig. 6. It categorizes questions into three main types: "context" (120 questions), “character” (40 questions), and “attribute” (140 questions). The “context” category includes subtypes such as relationships (22), activity (24), expertise (31), concrete (21), and abstract (22) questions. The “character” category comprises character/number questions (40). The “attribute” category encompasses presence (22), color/texture (24), number (24), shape (24), posture (24), and position (22) questions. This distribution highlights a comprehensive and balanced approach, ensuring a well-rounded benchmark with detailed subcomponents across various dimensions of question types.
8.3 Training Details
@@
8.4 Discussion 1: Can We Simply Introduce Negative Samples to Train MLLMs?
@@
1. If there is no text/number recognized or you do not make sure the text/number is correct, do not ask this type of question. If text is present, inquire about modifications to characters or numbers: Generate questions involving minor alterations, such as reversing the order of digits (modifying a number) or replacing a character. 2. If characters or numbers are present in the image, modify one letter of a word or one digit of a number, then ask if the modified character or number exists, and add a distractor sentence to make the question more challenging. Otherwise, do not ask this type of question. 3. Ask if certain objects exist: Replace the nouns of objects in the image with visually similar but nonexistent objects, then ask if the model identifies the object. 4. Ask if objects exist and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. 5. If objects have attributes such as names of people, colors of objects, counts of objects, or positions, ask modified questions about object attributes and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. 6. Ask global understanding questions: Understand the broader context, culture, emotions, and knowledge depicted in the image. Pose questions that differ slightly from true understanding but are very similar. 7. Ask global understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. 8. Ask local understanding questions: Based on the extracted relationships between objects, between people and objects, or between people, actions, or relative positions, pose local understanding challenges (negative samples). 9. Ask local understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. Requirements: 1. Generate 7-10 questions. 2. Ask directly, without adding any information such as "in the conversation," "mentioned in the conversation," or "in the image." 3. Answer directly, with explanation, succinctly, with responses not exceeding 30 characters. 4. Questions must be real and answerable. Answers must be correct, definitely,and not speculative. 5. Do not generate questions that the answer is "uncertain". 6. The format of the generated questions and answers is: ’conversations’: [’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, …, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’] STEP-3: Output JSON sample
@@
1. Character: (a) If no text/number is recognized or the text/number recognized is not sure, do not ask questions about text/number. If text exists, inquire about modifications to characters or numbers: Generate questions involving minor alterations, such as reversing the order of digits (modifying a number) or replacing a character. (b) If there are characters or numbers in the image, modify one letter or one digit, then ask if the modified character or number exists, and add a distractor sentence to make the question more challenging. Otherwise, do not ask this type of question. 2. Semantic: (a) Ask if certain objects exist: Replace the nouns of objects in the image with visually similar but nonexistent objects, then ask if the model identifies the object. (b) Ask if objects exist and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. (c) If objects have attributes such as names of people, colors of objects, counts of objects, or positions, ask modified questions about object attributes and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. 3. Understanding: (a) Ask global understanding questions: Understand the broader context, culture, emotions, and knowledge depicted in the image. Pose questions that differ slightly from true understanding but are very similar. (b) Ask global understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to make the question more challenging. (c) Ask local understanding questions: Based on the extracted relationships between objects, between people and objects, or between people, actions, or relative positions, pose local understanding challenges (negative samples). (d) Ask local understanding questions and add a distractor sentence: Add a distractor sentence to questions based on modifying characters or numbers to guide the model to make incorrect inferences or answers. Requirements: 1. Generate 2 character-type questions, 4 semantic-type questions, and 2 understanding-type questions. 2. Ask directly, without adding any information such as "in the conversation," "mentioned in the conversation," or "in the image." 3. Answer directly, with explanation, succinctly, with responses not exceeding 30 characters. 4. Questions must be real and answerable. Answers must be correct, definitely,and not speculative. 5. Do not generate questions that the answer is "uncertain". 6. The format of the generated questions and answers is: ’conversations’: [’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’, …, ’from’: ’human’, ’value’: ’Question’, ’from’: ’gpt’, ’value’: ’Answer’] STEP-3: Output JSON sample
@@
8.4.1 Training and Results of the Version 0.
@@
8.4.2 Training and Results of the Version 1.
@@
8.5 Discussion 2: What Constitutes an Effective Pair of Positive and Negative Samples for Training?
Having established the benefit of incorporating both positive and negative samples during training, we now delve into effective construction methods for such pairs. We investigate two distinct approaches: 1) Explicit Positive/negative Prompting. This method involves directly prompting the model to generate both positive and negative samples, with answers explicitly labeled as "yes" or "no." This approach offers straightforward control over the positive/negative balance in the training data (version 2). Confounding Answer Construction. This approach focuses on crafting questions that naturally lend themselves to multiple plausible answers, without explicitly forcing a "yes/no" response. This strategy aims to create a more nuanced and challenging training dataset, potentially enhancing the model’s ability to discern subtle differences in meaning and context (version 3). The GPT-4V prompts for version 2 and version 3 are shown in Fig. 13 and Fig. 14, respectively. To evaluate the effectiveness of these two versions of prompts, we generated 24k positive-negative sample pairs using each version’s prompt. We then incorporated these datasets into our training pipeline, maintaining consistent experimental settings with previous trials. Fig. 15 presents the training loss curves observed for both approaches, while Tab. 14 and Tab. 15 detail the corresponding performance results across various benchmarks. Our experimental results demonstrate that utilizing prompt version 3 to generate paired positive and negative samples yielded the strongest overall performance. Consequently, we adopted version 3 as our preferred method for generating training data in the main experiments.